Sultan Services has 1.1 million shares outstanding. It expects earnings at the end of the year of $ 5.00 million. Sultan pays out​ 60% of its earnings in total​ - 40% paid out as dividends and​ 20% used to repurchase shares. If​ Sultan's earnings are expected to grow by 7​% per​ year, these payout rates do not​ change, and​ Sultan's equity cost of capital is 9​%, what is​ Sultan's share​ price?

A. $ 81.82

B. $ 27.27

C. $ 54.54

D. $ 136.36
Answer #1

Expected dividend = earnings * payout ratio

= 5 * 0.6

= 3

Value of company = Expected dividend/cost of capital-growth rate

= 3/(0.09-0.07)

= 150 million

Price per share = value/number of shares

= 150/1.1

= 136.36

choose D)